OpenStudy (anonymous):

I have the Calculus BC final tomorrow. Any last minute study tips? I'm pretty much good to go, curriculum wise, but I'm scared out of my mind!

OpenStudy (anonymous):

Did you just delete something that you posted? I didn't quite catch that.

OpenStudy (anonymous):

Sample problem: \[\int\limits \frac{d \beta}{\left( \beta^2 + \xi^2\right)^{\frac{n}{2}}}\] :p

OpenStudy (anonymous):

0_0 Ok. Well. In that case I guess I should prepare myself to fail the test xD The beta thing throws me off... Are you certain that this is High school level Calc BC? Because I dont believe I've ever integrated something with so many variables in it! Hahaha

OpenStudy (anonymous):

Well you've done like: \[\int\limits \frac{dx}{(x^2+49)^{\frac{3}{2}}}\] right? Its the same thing but with beta exchanged for x and xi is a constant :P. The n/2 makes it more general and significantly harder, I was just messing really :P I'm in quantum mechanics right now and check out my hw:

OpenStudy (anonymous):

Sure, I've integrated that before. But I figured you were just messing. I genuinely attempted the problem though XD Way to pull at my nerves! haha. And quantum mechanics is pretty impressive. Good job. I wish I had the mental capacity to follow something like that.

OpenStudy (anonymous):

If you're interested I could probably work that problem out for you :P

OpenStudy (anonymous):

I would love that. But only if you wish. I would probably make the 3/2 a negative, so that f(x) could be in the numerator, then derive from there. Unless there is some sort of trig function formula that would work better. I don't think u substitution will work.

OpenStudy (anonymous):

Since 49 is a perfect square, im assuming there is some sort of formula that has slipped my mind.

OpenStudy (anonymous):

Well no need for formulas :P Rewrite it as: \[\int\limits \frac{dx}{(x^2+7^2)^{\frac{3}{2}}}\] And let: \[x = 7 \tan(\phi) \implies dx = 7 \sec^2(\phi)d \phi\] This gives us: \[\int\limits \frac{7 \sec^2(\phi) d \phi}{7^3(\tan^2(\phi)+1)^{\frac{3}{2}}}\] But: \[\cos^2(\phi)+\sin^2(\phi) = 1 \implies \frac{\cos^2(\phi)}{\cos^2(\phi)} + \frac{\sin^2(\phi)}{\cos^2(\phi)} = \frac{1}{\cos^2(\phi)} \] \[\implies 1 + \tan^2(\phi) = \sec^2(\phi)\] So we now have: \[\frac{1}{7^2} \int\limits \frac{\sec^2(\phi)d \phi}{(\sec^2(\phi))^{\frac{3}{2}}} = \frac{1}{49}\int\limits \frac{d \phi}{\sec(\phi)} = \frac{1}{49} \int\limits \cos(\phi) d \phi = \frac{1}{49}\sin(\phi)+C\]

OpenStudy (anonymous):

That problem>My understanding. xD But thank you! It is very impressive. I wish you could take my final for me tomorrow, you seem very well prepared. ;)

OpenStudy (anonymous):

However:  So we have that: \[\sin(\phi) = \frac{\textrm{opposite}}{\textrm{hypotenuse}} = \frac{x}{\sqrt{x^2+49}}\] So our final answer is: \[\frac{x}{49} \frac{1}{\sqrt{x^2+49}} + C\] You haven't done trig subs? :O

OpenStudy (anonymous):

Not at such a magnitude, no. We have, but more simplistically, I believe. Simple, precalc-ish trig subs. Nothing like that.

OpenStudy (anonymous):

Oh :x my bad then lol However, if you have any other problems you don't know fire away!

OpenStudy (anonymous):

Thanks! I wont hesitate to ask. As you can see, I'm pretty ill-prepared. Lots of little mistakes and slip ups that need refining. Thank you for your help.

OpenStudy (anonymous):

Not a problem, I get to keep my calc skills sharp as well :P